Let Z be the standard normal variable. Find z if z satisfies the given value. (Round your answer to two decimal places.) P(−z < Z < z) = 0.9802
Answer #1

solution
P(-z < Z < z) =0.9802

P(Z < z) - P(Z < -z) = 0.9802

2 P(Z < z) - 1 =0.9802

2 P(Z < z) = 1 + 0.9802. = 1.9802

P(Z < z) = 1.9802 / 2 = 0.9901

P(Z < 2.33) = 0.9901

z  ± 2.33 using z table

z= -2.33 , z= + 2.33